allcm(1) teTeX allcm(1) NAME allcm−forcethemostimportantComputer−Modern−fontstobecalculated SYNOPSIS allcm DESCRIPTION allcmforcesalargenumberofComputerModernFontstobecalculatedaspixelfiles.Thisis donethroughrunningdvips(1)overacertaintest−file.Therefore,thefontsarecreatedintheresolutionneededbydvips. allecdoesnotrecalculateexistingfonts(aslongastheMetafontmodedoesnotchange). SEEALSO allec(1),allneeded(1),fontimport(1),fontexport(1). AUTHOR ThomasEsser(te@informatik.uni−hannover.de) teTeX 11/94 1